the equation of the diagonal through the origin of the quadrilateral formed by x=0 , y=0 , x+y=1 and 6 X + Y = 3 is given by

Grade:11

1 Answers

Vikas TU
7 years ago
After plotting all the lines the diagonal becomes the line joining the origin to the intersection points of the eqns.
x + y = 1 and 6x + y = 3 respective;y.
The intersection points aree: (2/5,3/5)
Hence the slope, m = 3/2
AN dthe eqn. of the straight line is: => y = 3x/2